OpenStudy (anonymous):

amy draes one card from one bag and another card from another bag the bags contain 26 letters and 10 numerals, whats is the probability that she draws 2 letters

OpenStudy (anonymous):

So both bags contain the same number of cards? If that's the case, then bags 1 and 2 have 36 total cards each (assuming each card has either one number or one letter on it). The probability of drawing a card with a number on it from either bag is \(\dfrac{10}{36}\) because 10 of the 36 cards have numbers. This means the probability of drawing a number from both bags is \(\dfrac{10}{36}\times\dfrac{10}{36}\).